OpenStudy (anonymous):

(csc-1)(csc+1)=cot^2 csc^2-1=cot^2 cot^2=cot^2

OpenStudy (anonymous):

@Michele_Laino is this correct?

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

Awesome, thank you, how much longer will you be on?

OpenStudy (michele_laino):

I stay here, for at least 2 hours

OpenStudy (anonymous):

\[\cos^2\theta(1+\tan ^2\theta)=1\] \[\cos^2\theta(\sec^2\theta)=1\] \[(\cos^2\theta/1)(1/\cos^2\theta)=1\] 1=1

OpenStudy (anonymous):

Yeah! That's correct!

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

\[\sec \theta - \tan \theta= \cos \theta/1+\sin \theta\] \[(1/\cos)-(\sin/\cos)= \cos/1+\sin\] Then would the next line by 1-sin

OpenStudy (michele_laino):

yes! it would be: \[\frac{{1 - \sin }}{{\cos }}\]

OpenStudy (anonymous):

the cos wouldn't cancel?

OpenStudy (michele_laino):

no, since the right side is: \[\frac{{\cos }}{{1 + \sin }}\]

OpenStudy (michele_laino):

we can use this identity: \[\cos = \sqrt {1 - {{\sin }^2}} = \sqrt {1 - \sin } \sqrt {1 + \sin } \]

OpenStudy (anonymous):

Is there another identity that can be used? Our teacher doesnt have that one as an identity we can use

OpenStudy (michele_laino):

for example, the left side, will be like below: \[\large \frac{{1 - \sin }}{{\cos }} = \frac{{1 - \sin }}{{\sqrt {1 - \sin } \times \sqrt {1 + \sin } }} = \frac{{\sqrt {1 - \sin } \times \sqrt {1 - \sin } }}{{\sqrt {1 - \sin } \times \sqrt {1 + \sin } }} = ...?\]

OpenStudy (michele_laino):

please, wait, I'm searching for another method

OpenStudy (michele_laino):

here it is another method: \[\frac{{1 - \sin }}{{\cos }} = \frac{{1 - \sin }}{{\cos }} \cdot \frac{{\cos }}{{\cos }} = \frac{{\cos \left( {1 - \sin } \right)}}{{{{\cos }^2}}} = \frac{{\cos \left( {1 - \sin } \right)}}{{1 - {{\sin }^2}}} = \frac{{\cos \left( {1 - \sin } \right)}}{{\left( {1 - \sin } \right)\left( {1 + \sin } \right)}} = ...?\]

OpenStudy (michele_laino):

\[\begin{gathered} \frac{{1 - \sin }}{{\cos }} = \frac{{1 - \sin }}{{\cos }} \cdot \frac{{\cos }}{{\cos }} = \frac{{\cos \left( {1 - \sin } \right)}}{{{{\cos }^2}}} = \frac{{\cos \left( {1 - \sin } \right)}}{{1 - {{\sin }^2}}} = \hfill \\ = \frac{{\cos \left( {1 - \sin } \right)}}{{\left( {1 - \sin } \right)\left( {1 + \sin } \right)}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

namely I have multiplied and divided, the left side by cos(theta)

OpenStudy (michele_laino):

so we can write: 

OpenStudy (anonymous):

Why didn't distribute the cos?

OpenStudy (michele_laino):

because we have to solve an identity, so we have to start from the left side, and using some computation, we have to transform that left side into the right side

OpenStudy (michele_laino):

or, similarly starting from the right side, we have to transform that right side into the left side

OpenStudy (anonymous):

And what happened to the cos^2

OpenStudy (michele_laino):

I can rewrite cos^2 as below: cos^2 = 1-sin^2 =(1-sin)(1+sin) namely I have used the fundamental identity

OpenStudy (anonymous):

why wouldnt it be two (1-sin)?

OpenStudy (michele_laino):

they cancel each other, since one factor (1-sin) is at the numerator, and the second one is at denominator

OpenStudy (michele_laino):

there I have applied this algebraic identity: a^2-b^2= (a-b)(a+b)

OpenStudy (michele_laino):

where a^2= 1, and b^2 = (sin)^2, so a= 1 and b = sin(theta)

OpenStudy (anonymous):

Okay I understand

OpenStudy (anonymous):

\[9\sec^2\theta-5\tan^2\theta=5+4\sec^2\theta\]

OpenStudy (michele_laino):

here, we can start from the left side, and we can write this: \[\Large \begin{gathered} 9{\left( {\sec \theta } \right)^2} - 5{\left( {\tan \theta } \right)^2} = \hfill \\ \\ =9\frac{1}{{{{\left( {\cos \theta } \right)}^2}}} - 5\frac{{{{\left( {\sin \theta } \right)}^2}}}{{{{\left( {\cos \theta } \right)}^2}}} = \frac{{9 - 5{{\left( {\sin \theta } \right)}^2}}}{{{{\left( {\cos \theta } \right)}^2}}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

I dont know what to do next

OpenStudy (michele_laino):

hint: \[\Large \begin{gathered} \frac{{9 - 5{{\left( {\sin \theta } \right)}^2}}}{{{{\left( {\cos \theta } \right)}^2}}} = \frac{{4 + 5 - 5{{\left( {\sin \theta } \right)}^2}}}{{{{\left( {\cos \theta } \right)}^2}}} = \hfill \\ \hfill \\ = \frac{{4 + 5\left( {1 - {{\left( {\sin \theta } \right)}^2}} \right)}}{{{{\left( {\cos \theta } \right)}^2}}} = \frac{{4 + 5{{\left( {\cos \theta } \right)}^2}}}{{{{\left( {\cos \theta } \right)}^2}}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Where did the -5 go?

OpenStudy (michele_laino):

I have factored out 5, so I could write: \[\Large 5 - 5{\left( {\sin \theta } \right)^2} = 5\left( {1 - {{\left( {\sin \theta } \right)}^2}} \right)\]

OpenStudy (anonymous):

I understand that now, then what happens next?

OpenStudy (michele_laino):

we have the subsequent steps: \[\Large \frac{{4 + 5{{\left( {\cos \theta } \right)}^2}}}{{{{\left( {\cos \theta } \right)}^2}}} = \frac{4}{{{{\left( {\cos \theta } \right)}^2}}} + \frac{{5{{\left( {\cos \theta } \right)}^2}}}{{{{\left( {\cos \theta } \right)}^2}}} = \frac{4}{{{{\left( {\cos \theta } \right)}^2}}} + 5\]

OpenStudy (anonymous):

Got that part now

OpenStudy (anonymous):

Do we change the cos next?

OpenStudy (michele_laino):

we can use this identity: \[\Large \frac{1}{{{{\left( {\cos \theta } \right)}^2}}} = {\left( {\sec \theta } \right)^2}\]

OpenStudy (michele_laino):

so we get: \[\Large \frac{4}{{{{\left( {\cos \theta } \right)}^2}}} + 5 = 4{\left( {\sec \theta } \right)^2} + 5\]

OpenStudy (anonymous):

Got it

OpenStudy (anonymous):

Sorry its so sloppy, I didnt feel like typing it out

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

your method is right!

OpenStudy (anonymous):

Would I multiple the cot to the 1, and the 1 on top of cos

OpenStudy (michele_laino):

here is the next step: \[\Large \frac{{1 + \tan }}{{1 - \tan }} = \frac{{1 + \frac{1}{{\cot }}}}{{1 - \frac{1}{{\cot }}}} = \frac{{\frac{{\cot + 1}}{{\cot }}}}{{\frac{{\cot - 1}}{{\cot }}}} = \frac{{\cot + 1}}{{\cot }} \times \frac{{\cot }}{{\cot - 1}} = ...?\]

OpenStudy (michele_laino):

\[\Large \begin{gathered} \frac{{1 + \tan }}{{1 - \tan }} = \frac{{1 + \frac{1}{{\cot }}}}{{1 - \frac{1}{{\cot }}}} = \frac{{\frac{{\cot + 1}}{{\cot }}}}{{\frac{{\cot - 1}}{{\cot }}}} = \hfill \\ \hfill \\ = \frac{{\cot + 1}}{{\cot }} \times \frac{{\cot }}{{\cot - 1}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Did you multiple cot to something to get cot+1

OpenStudy (michele_laino):

no, since I have computed this: \[\Large 1 + \tan = 1 + \frac{1}{{\cot }} = \frac{1}{1} + \frac{1}{{\cot }} = \frac{{1 \cdot \cot + 1}}{{\cot }} = \frac{{1 + \cot }}{{\cot }}\]

OpenStudy (michele_laino):

namely I have summed the 2 fraction 1 and 1/cot each other

OpenStudy (michele_laino):

fractions*

OpenStudy (michele_laino):

similarly for: 1-tan

OpenStudy (anonymous):

I still dont understand sorry

OpenStudy (michele_laino):

no worries, I retry, please wait a moment

OpenStudy (michele_laino):

we have to sum this quantities: 1=1/1 and 1/cot(theta) Now the least common multiple between 1 and cot(theta) is: cot(theta)

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

so we have: \[\Large 1 + \tan = 1 + \frac{1}{{\cot }} = \frac{1}{1} + \frac{1}{{\cot }} = \frac{{1 \cdot \cot + 1}}{{\cot }}\]

OpenStudy (michele_laino):

\[\Large \frac{1}{1} + \frac{1}{{\cot }} = \frac{{1 \cdot \cot + 1}}{{\cot }}\]

OpenStudy (michele_laino):

for example if I have to sum these fractions: 1 and 1/5, then the least common multiple is 5, so Ican write: \[\Large 1 + \frac{1}{5} = \frac{1}{1} + \frac{1}{5} = \frac{{1 \times 5 + 1}}{5}\]

OpenStudy (michele_laino):

oops..so I can write...

OpenStudy (anonymous):

Ohhh I get it now

OpenStudy (anonymous):

do you have to go?

OpenStudy (michele_laino):

no, I stay here

OpenStudy (anonymous):

okay, I'm going to start a new post because this one is getting too long

OpenStudy (michele_laino):

ok!